OpenStudy (anonymous):

if k objects out of a set of n elements are placed into a fixed order, how many unique arrangements of those objects are there?

OpenStudy (anonymous):

@amistre64 im soo lost...

OpenStudy (amistre64):

since we cant express it as some nice little number setup, we have to define it by a formula

OpenStudy (amistre64):

i defined the formula for it the last time ... but after each time i stated it you were still "confused" n!/(n-k)! is the formula

OpenStudy (anonymous):

ok.. i get that there's a formula, but i dont get how you decide which numbers to use.

OpenStudy (amistre64):

a,b,c,d,e,f if we take any 3 from the set and place them in a unique order; we get 60 different subsets. 5!/(5-3)! = 5*4*3 = 20*3 = 60

OpenStudy (amistre64):

if we take k elements from a set of n elements ,,,, we can form: n!/(n-k)! unique arrangements

OpenStudy (anonymous):

where do you get 5 from..

OpenStudy (amistre64):

hmm, i thought i only used 5 letters :) spose we ignore the f

OpenStudy (amistre64):

the number of elements of a set doesnt care with the elements are called; and by definition the elements that make a set are uniquly defined ... no duplicates

OpenStudy (anonymous):

im sorry.. i am still really confused... im not just saying that so that you give me the answer, but i am homeschooled and new to this section, and i dont get the whole "!" thing or the equation and how it all works..

OpenStudy (amistre64):

the set {a,b,c,d,e} has 5 elements the number of ways to use 3 of those elements .....

OpenStudy (amistre64):

! is a notation ... something that helps to shorten the way we write a number is all

OpenStudy (amistre64):

do you recall the notation from multiple additions? say; 3+3+3+3+3+3+3 = 7(3) do you recall the notation from multiple multiplications? say; 3*3*3*3*3*3*3 = 3^7 the notation ! expresses a different way of calculating a certain number

OpenStudy (anonymous):

ya

OpenStudy (amistre64):

3! is a way of expressing the number: 3*2*1 = 6 7! is a way of expressing the number: 7*6*5*4*3*2*1 = 5040 its just a notation for the way of expressing the value of the product of successive values starting at a certain number and going to 1

OpenStudy (anonymous):

okk

OpenStudy (anonymous):

so if it doesnt say a number, it just says a variable, how do you know what number?

OpenStudy (amistre64):

the question tells you to use n and k .....

OpenStudy (amistre64):

the formula for calculating the number of permutations of a set of "n" elements, by using "k" at a time amounts to: n!/(n-k)!

OpenStudy (anonymous):

?ok.. but how do you solve it

OpenStudy (amistre64):

by using the definition of a factorial notation ....

OpenStudy (amistre64):

lets use the example of 5 elements ordered 3 at a time ... what is 5! ?

OpenStudy (amistre64):

3! is a way of expressing the number: 3*2*1 = 6 7! is a way of expressing the number: 7*6*5*4*3*2*1 = 5040 its just a notation for the way of expressing the value of the product of successive values starting at a certain number and going to 1 so what is the value of 5! ?

OpenStudy (anonymous):

120

OpenStudy (anonymous):

but there are no numbers with k and n

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

anyone......???